# Eccentric Mergers in AGN Discs: Influence of the Supermassive Black-Hole on Three-body Interactions

### v1

@@ Abstract
There are indications that stellar-origin black holes (BHs) are efficiently paired up in binary black holes (BBHs) in Active Galactic Nuclei (AGN) disc environments, which can undergo interactions with single BHs in the disc. Such binary-single interactions can potentially lead to an exceptionally high fraction of gravitational-wave mergers with measurable eccentricity in LIGO/Virgo/KAGRA. We here take the next important step in this line of studies, by performing post-Newtonian N-body simulations between migrating BBHs and single BHs set in an AGN disc-like configuration with a consistent inclusion of the central supermassive black hole (SMBH) in the equations of motion. With this setup, we study how the fraction of eccentric mergers varies in terms of the initial size of the BBH semi-major axis relative to the Hill sphere, as well as how it depends on the angle between the BBH and the incoming single BH. We find that the fraction of eccentric mergers is still relatively large, even when the interactions are notably influenced by the gravitational field of the nearby SMBH. However, the fraction as a function of the BBH semi-major axis does not follow a smooth functional shape, but instead shows strongly varying features that originate from the underlying phase-space structure. The phase-space further reveals that many of the eccentric mergers are formed through prompt scatterings. Finally, we present the first analytical solution to how the presence of an SMBH in terms of its Hill sphere affects the probability for forming eccentric BBH mergers through chaotic three-body interactions.
@@
stars: kinematics and dynamics – galaxies: active – accretion, accretion discs – galaxies: nuclei – stars: black holes – gravitational waves
@@ 1 Introduction
The observations of gravitational-wave (GW) sources with LIGO/Virgo/KAGRA (LVK) is ongoing and have to date revealed sources of both binary black holes (BBHs) (Abbott et al., 2016b, c, a, 2017a, 2017b; Venumadhav et al., 2020a, b), binary neutron stars (BNSs) (Abbott et al., 2017c), as well as possible black hole-neutrons star mergers (Abbott et al., 2023; Hamers et al., 2021; Vynatheya & Hamers, 2022). However, how and where these sources form in our Universe are still major unsolved problems. Many formation channels have been proposed, including field binaries (Dominik et al., 2012, 2013, 2015; Kinugawa et al., 2014; Belczynski et al., 2016b; Belczynski et al., 2016a; Silsbee & Tremaine, 2017; Murguia-Berthier et al., 2017; Rodriguez & Antonini, 2018; Schrøder et al., 2018), dense stellar clusters (Portegies Zwart & McMillan, 2000; Lee et al., 2010; Banerjee et al., 2010; Tanikawa, 2013; Bae et al., 2014; Rodriguez et al., 2015; Rodriguez et al., 2016a, b, b; Askar et al., 2017; Park et al., 2017; Samsing, 2018; D’Orazio & Samsing, 2018; Samsing & D’Orazio, 2018; Samsing et al., 2020; Rozner & Perets, 2022), active galactic nuclei (AGN) accretion discs (Bartos et al., 2017; Stone et al., 2017; McKernan et al., 2018; Tagawa et al., 2019; Rozner et al., 2023), galactic nuclei (GN) (O’Leary et al., 2009; Hong & Lee, 2015; VanLandingham et al., 2016; Antonini & Rasio, 2016; Stephan et al., 2016; Hoang et al., 2018; Hamers et al., 2018), very massive stellar mergers (Loeb, 2016; Woosley, 2016; Janiuk et al., 2017; D’Orazio & Loeb, 2018), and single-single GW captures of primordial black holes (Bird et al., 2016; Cholis et al., 2016; Sasaki et al., 2016; Carr et al., 2016). To disentangle these different scenarios using GWs, we have to understand how the observable parameters differ between the different channels. For example, studies indicate that one can distinguish dynamically induced mergers from isolated binaries, by analyzing the the relative spin orientation of the merging black-holes (BHs) (Rodriguez et al., 2016c), the orbital eccentricity at some reference GW frequency (Gültekin et al., 2006; Samsing et al., 2014; Samsing & Ramirez-Ruiz, 2017; Samsing & Ilan, 2018; Samsing et al., 2018a; Samsing, 2018; Samsing et al., 2018b; Samsing & D’Orazio, 2018; Rodriguez et al., 2018; Zevin et al., 2019; Samsing et al., 2019b; Samsing et al., 2020), as well as the mass spectrum (Zevin et al., 2017). The environment in which the BBH was formed and merged can also be imprinted in the GW form, showing up as e.g. a GW phase-shift (e.g. Inayoshi et al., 2017; D’Orazio & Loeb, 2020, Samsing & Hendriks (in prep.)). Other probes of formation include e.g. stellar tidal disruptions (e.g. Samsing et al., 2019a; Lopez et al., 2019; Kremer et al., 2019b). From these studies it generally follows that dynamically formed mergers tend to have mass ratios near one (e.g. Rodriguez et al., 2018), random spin orientations(e.g. Rodriguez et al., 2016c), as well as a non-negligible fraction of sources with measurable eccentricity in LISA (Samsing & D’Orazio, 2018; D’Orazio & Samsing, 2018; Kremer et al., 2019a), DECIGO/Tian-Qin (e.g. Chen & Amaro-Seoane, 2017; Samsing et al., 2020), and LVK (Samsing, 2018). This is in contrast to isolated binary mergers, which are more likely to have a non-random spin distribution (e.g. Kalogera, 2000; Zaldarriaga et al., 2018; Hotokezaka & Piran, 2017; Piran & Piran, 2020), larger mass ratios, and eccentricity $\approx 0$ near LISA, DECIGO/Tian-Qin, and LVK.
However, some of the observed GW mergers start to challenge these classical formation channels. For example, GW190521 (Abbott et al., 2020b, c) seems to have masses that are either above the so called BH mass gap (e.g Fishbach & Holz, 2020), or one above and one below (Fishbach et al., 2020), which certainly was not generally expected from any models prior to this observation. In addition, GW190521 showed indication of non-zero eccentricity at 10 Hz, and a possible corresponding electro-magnetic (EM) counterpart (e.g. Graham et al., 2020). These features have led to the proposal that GW190521 could have formed in an AGN disc, where BHs can grow to high masses through gas accretion (Jermyn et al., 2021; Cantiello et al., 2021) or repeated mergers (Tagawa et al., 2021). They can also encounter each other in disc-like configurations that have been shown to produce up to two orders of magnitude more eccentric mergers compared to isotropic environments (Samsing et al., 2022) (from hereon referred to as S22), as well as interact with the surrounding gas to create possible EM counterparts (e.g. Graham et al., 2023). This observation, the underlying rich physics that brings together stellar-origin black holes, supermassive black holes, gas, dynamics and disc-accretion physics, have opened up a wealth of studies on how BBHs might form and merge in such AGN-disc environments (Tagawa et al., 2020). The general picture for this formation channel, is that BHs are either captured by the disc (e.g Bartos et al., 2017; Fabj et al., 2020; MacLeod & Lin, 2020; Nasim et al., 2023; Generozov & Perets, 2023; Wang et al., 2023), or formed in the disc as a result of star formation in the outskirts (Stone et al., 2017; Artale et al., 2019; Cantiello et al., 2021; Jermyn et al., 2022). We envision this population of black-holes to then migrate through the disc during which they may pair up to form binaries. These binaries are either brought to merge in the disc through a combination of gas drag and GW radiation (Tagawa et al., 2020), or surviving inside the disc for them to later encounter either other BBHs or single BHs (Tagawa et al., 2020; Samsing et al., 2022). In S22 it was argued that 3-body interactions, i.e. interactions between assembled BBHs and single BHs in the disc dominate the formation of BBH mergers, as each interaction hardens the BBH, possibly increase its eccentricity, and thereby bring the BBHs to merge on a short timescale. In S22, the first AGN-disc like 3-body simulations were performed with the inclusion of Post-Newtonian (PN) radiation terms (e.g. Blanchet, 2006, 2014), which showed that such interactions lead to an exceptionally high fraction of eccentric mergers. Other studies focusing on the formation process of the BBH itself under the influence of an SMBH, through processes known as Jacobi captures have also found that these BHs tend to pair up and possibly merge with very high eccentricity (Boekholt et al., 2023). Recent work (Trani et al., 2023) has performed complementary studies on how BHs interact in stellar-remnant discs environments that do not necessarily move on Keplerian orbits as expected in AGN-discs, and find that eccentric mergers are formed in large fraction. Recent studies have started to explore the effect of a gaseous disc in the problem of BH pairing with the inclusion of gas (Li & Lai, 2022; Li & Lai, 2023b; Li & Lai, 2023a; DeLaurentiis et al., 2023; Whitehead et al., 2023; Rozner et al., 2023; Rowan et al., 2023, 2024). While these studies are encouraging, the problem of gas friction exerted by the accretion disc when modeling the PN 3-body scattering problem in AGNs is a rather a challenging task. Therefore, in this work we neglect the role of gas and focus on the other prominent environmental effect in AGNs: the tidal field exerted by the SMBH.
@@
[…]
@@ 2 Three-body Interactions in a Tidal Field
Figure 1: Illustration of binary-single interaction propagating from left to right resulting in a merger between two of the three objects while the remaining single is still bound (see Sec. 2.1). The system goes through several intermediate states (IMS) before GW capture of two of the three objects. For illustrative purposes, we include horizontal dashed lines to indicate the size of a Hill sphere ($R_{\rm H}$) in case the interaction takes place in the presence of a tidal field. The interaction undergoes a tidal breakup if it extends beyond this sphere. The tidal boundary limits the number of interactions $\mathcal{N}$ by breaking up the system before natural completion, which therefore leads to a reduction in the probability for 3-body and eccentric mergers. The derived correction factor is shown in Eq. 22.
The three-body scattering problem has been studied in great detail with different physical motivations, from the context of classical cluster evolution (Heggie, 1975) to including PN correction terms to follow the formation of GW driven mergers of compact objects (Samsing et al., 2014; Samsing, 2018). While several theoretical studies have been put forward for describing especially the fraction of eccentric merges forming in chaotic 3-body interactions (e.g. Samsing et al., 2014; Rodriguez et al., 2016a; Samsing, 2018), only a few have extended this theory to include how a nearby perturber, such as an SMBH, affect the range, fraction, and nature of outcomes (e.g. Trani et al., 2019; Ginat & Perets, 2021; Trani et al., 2023; Rom et al., 2023). In S22 it was shown analytically and numerically that co-planar scatterings, such as those that might take place in AGN-disc environments, naturally give rise to an exceptionally high fraction of eccentric GW mergers due to the controlled geometric setup, neglecting however the effect of the SMBH. Below we start by reviewing the relevant 3-body scattering theory, after which we derive how the presence of a nearby object to leading order has an affect on the fraction of eccentric mergers forming during the chaotic state of the 3-body system. This not only adds an important component to this theoretical framework, but also sets the stage and provides motivation for full PN scatterings with an SMBH that we present later in the paper.
@@ 2.1 Formation of eccentric mergers
A 3-body merger is always highly eccentric at formation, but it might be that it forms with a GW peak frequency that is below the observable band of LVK (Samsing et al., 2014). In this case, one has to evolve the binary until it reaches the band through the emission of GWs; however, at this point the eccentricity might be too low to be measured. Therefore, 3-body mergers are not necessarily representative of the fraction of observable highly eccentric mergers in e.g. LVK as discussed in Samsing (2018). For an IMS merger to clearly show up as an eccentric merger, its GW peak frequency at formation has to be near the observable band. For a highly eccentric binary with a pericenter distance $r_{f}$, the corresponding GW peak frequency can be approximated by
@@
it consequently shows up eccentric at GW frequency that is higher than $f_{\rm gw}$. With this critical distance one can now easily estimate the probability for the outcome in a similar way to what we did in Eq. 8, which is $\approx\mathcal{N}\times\sqrt{2r_{f}/a_{0}}$. In the more general case for the IMS binary to appear at GW frequency $f_{\rm gw}$ with an eccentricity $e_{f}$, the probability should instead by based on the critical pericenter distance,
@@
as further explained in S22. The question is how do these general results, which also have shown excellent agreement with numerical simulations (S22), change when the chaotic three-body interactions take place in a tidal field from a nearby perturber, such as an SMBH. Below we estimate how this addition affects the outcomes.

### v2

@@ Abstract
There are indications that stellar-origin black holes (BHs) are efficiently paired up in binary black holes (BBHs) in Active Galactic Nuclei (AGN) disc environments, which can undergo interactions with single BHs in the disc. Such binary-single interactions can potentially lead to an exceptionally high fraction of gravitational-wave mergers with measurable eccentricity in LIGO/Virgo/KAGRA. We here take the next important step in this line of studies by performing post-Newtonian N-body simulations between migrating BBHs and single BHs set in an AGN disc-like configuration, with a consistent inclusion of the central supermassive black hole (SMBH) in the equations of motion. With this setup, we study how the fraction of eccentric mergers varies in terms of the initial size of the BBH semi-major axis relative to the Hill sphere, as well as how it depends on the angle between the BBH and the incoming single BH. We find that the fraction of eccentric mergers is still relatively large, even when the interactions are notably influenced by the gravitational field of the nearby SMBH. However, the fraction as a function of the BBH semi-major axis does not follow a smooth functional shape, but instead shows strongly varying features that originate from the underlying phase-space structure. The phase-space further reveals that many of the eccentric mergers are formed through prompt scatterings. Finally, we present the first analytical solution to how the presence of an SMBH in terms of its Hill sphere affects the probability for forming eccentric BBH mergers through chaotic three-body interactions.
@@
galaxies: active – galaxies: kinematics and dynamics – stars: black holes – gravitational waves
@@ 1 Introduction
The observations of gravitational-wave (GW) sources with LIGO/Virgo/KAGRA (LVK) are ongoing and have to date revealed sources of both binary black holes (BBHs) (Abbott et al., 2016b, c, a, 2017a, 2017b; Venumadhav et al., 2020a, b), binary neutron stars (BNSs) (Abbott et al., 2017c), as well as possible black-hole neutrons star mergers (Abbott et al., 2023; Hamers et al., 2021; Vynatheya & Hamers, 2022). However, how and where these sources form in our Universe remain major unsolved problems. Many formation channels have been proposed, including field binaries (Dominik et al., 2012, 2013, 2015; Kinugawa et al., 2014; Belczynski et al., 2016b; Belczynski et al., 2016a; Silsbee & Tremaine, 2017; Murguia-Berthier et al., 2017; Rodriguez & Antonini, 2018; Schrøder et al., 2018), dense stellar clusters (Portegies Zwart & McMillan, 2000; Lee et al., 2010; Banerjee et al., 2010; Tanikawa, 2013; Bae et al., 2014; Rodriguez et al., 2015; Rodriguez et al., 2016a, b, b; Askar et al., 2017; Park et al., 2017; Samsing, 2018; D’Orazio & Samsing, 2018a; Samsing & D’Orazio, 2018; Samsing et al., 2020; Rozner & Perets, 2022), active galactic nuclei (AGN) accretion discs (Bartos et al., 2017; Stone et al., 2017; McKernan et al., 2018; Tagawa et al., 2019; Rozner et al., 2023), galactic nuclei (GN) (O’Leary et al., 2009; Hong & Lee, 2015; VanLandingham et al., 2016; Antonini & Rasio, 2016; Stephan et al., 2016; Hoang et al., 2018; Hamers et al., 2018), very massive stellar mergers (Loeb, 2016; Woosley, 2016; Janiuk et al., 2017; D’Orazio & Loeb, 2018), and single-single GW captures of primordial black holes (Bird et al., 2016; Cholis et al., 2016; Sasaki et al., 2016; Carr et al., 2016). To disentangle these different scenarios using GWs, we have to understand how the observable parameters differ between the different channels. For example, studies indicate that one can distinguish dynamically induced mergers from isolated binaries, by analyzing the relative spin orientation of the merging black-holes (BHs) (Rodriguez et al., 2016c), the orbital eccentricity at some reference GW frequency (Gültekin et al., 2006; Samsing et al., 2014; Samsing & Ramirez-Ruiz, 2017; Samsing & Ilan, 2018; Samsing et al., 2018a; Samsing, 2018; Samsing et al., 2018b; Samsing & D’Orazio, 2018; Rodriguez et al., 2018; Zevin et al., 2019; Samsing et al., 2019b; Samsing et al., 2020), as well as the mass spectrum (Zevin et al., 2017). The environment in which the BBH was formed and merged can also be imprinted in the GW form, showing up as e.g. a GW phase-shift (e.g. Inayoshi et al., 2017; D’Orazio & Loeb, 2020; Samsing et al., 2024; Hendriks et al., 2024). Other probes of formation include e.g. stellar tidal disruptions (e.g. Samsing et al., 2019a; Lopez et al., 2019; Kremer et al., 2019b). From these studies it generally follows that dynamically formed mergers tend to have mass ratios near one (e.g. Rodriguez et al., 2018), random spin orientations (e.g. Rodriguez et al., 2016c), as well as a non-negligible fraction of sources with measurable eccentricity in LISA (Samsing & D’Orazio, 2018; D’Orazio & Samsing, 2018a; Kremer et al., 2019a), DECIGO/Tian-Qin (e.g. Chen & Amaro-Seoane, 2017; Samsing et al., 2020), and LVK (Samsing, 2018). This is in contrast to isolated binary mergers, which are more likely to have a non-random spin distribution (e.g. Kalogera, 2000; Zaldarriaga et al., 2018; Hotokezaka & Piran, 2017; Piran & Piran, 2020), larger mass ratios, and eccentricity $\approx 0$ near LISA, DECIGO/Tian-Qin, and LVK.
However, some of the observed GW mergers start to challenge these classical formation channels. For example, GW190521 (Abbott et al., 2020b, c) seems to have masses that are either above the so called BH mass gap (e.g. Fishbach & Holz, 2020), or one above and one below (Fishbach et al., 2020; Nitz & Capano, 2021), which certainly was not generally expected from any models prior to this observation. In addition, GW190521 suggests better consistency with an eccentric waveform with $e\geq 0.1$ at 10 Hz, rather than a quasi-circular one (e.g. Romero-Shaw et al., 2020, 2022; Gayathri et al., 2022), and a possible corresponding electro-magnetic (EM) counterpart (e.g. Graham et al., 2020). These features have led to the proposal that GW190521 could have formed in an AGN disc, where BHs can grow to high masses through gas accretion (e.g. McKernan et al., 2012; Gilbaum & Stone, 2022) or repeated mergers (Tagawa et al., 2021). They can also encounter each other in disc-like configurations that have been shown to produce up to two orders of magnitude more eccentric mergers compared to isotropic environments (Samsing et al., 2022) (from hereon referred to as S22), as well as interact with the surrounding gas to create possible EM counterparts (e.g. Graham et al., 2023).
This observation, the underlying rich physics that brings together stellar-origin black holes, supermassive black holes, gas, dynamics and disc-accretion physics, have opened up a wealth of studies on how BBHs might form and merge in such AGN-disc environments (Tagawa et al., 2020). The general picture for this formation channel is that BHs can be formed through different mechanisms. They can be captured by the disc (e.g. Bartos et al., 2017; Fabj et al., 2020; MacLeod & Lin, 2020; Nasim et al., 2023; Generozov & Perets, 2023; Wang et al., 2024), form in the disc as a result of star formation in the outskirts (Stone et al., 2017; Artale et al., 2019), or they can be the result of core collapse of stars that undergo rapid mass accretion in high-density regions of the accretion disc (Jermyn et al., 2021; Cantiello et al., 2021; Dittmann et al., 2021; Jermyn et al., 2022; Dittmann et al., 2023b).
We envision this population of black-holes to then migrate through the disc during which they may pair up to form binaries. These binaries are either brought to merge in the disc through a combination of gas drag and GW radiation (Tagawa et al., 2020; Li & Lai, 2022; Dempsey et al., 2022; Li & Lai, 2023b; Li & Lai, 2023a; Dittmann et al., 2024; Calcino et al., 2024), or surviving inside the disc for them to later encounter either other BBHs or single BHs (Tagawa et al., 2020; Samsing et al., 2022). In S22 it was argued that 3-body interactions, i.e. interactions between assembled BBHs and single BHs in the disc dominate the formation of BBH mergers, as each interaction hardens the BBH, possibly increase its eccentricity, and thereby bring the BBHs to merge on a short timescale. In S22, the first AGN-disc like 3-body simulations were performed with the inclusion of Post-Newtonian (PN) radiation terms (e.g. Blanchet, 2006, 2014), which showed that such interactions lead to an exceptionally high fraction of eccentric mergers.
Other studies focusing on the formation process of the BBH itself under the influence of an SMBH, through processes known as Jacobi captures, have also found that these BHs tend to pair up and possibly merge with very high eccentricity (Boekholt et al., 2023). Recent work (Trani et al., 2023) has performed complementary studies on how BHs interact in stellar-remnant discs environments that do not necessarily move on Keplerian orbits as expected in AGN-discs, and find that eccentric mergers are formed in large fraction. Recent studies have started to explore the effect of a gaseous disc in the problem of BH pairing with the inclusion of gas (Li et al., 2023; DeLaurentiis et al., 2023; Whitehead et al., 2024; Rozner et al., 2023; Rowan et al., 2023, 2024). While these studies are encouraging, the problem of gas friction exerted by the accretion disc when modeling the PN 3-body scattering problem in AGNs is a rather challenging task. Therefore, in this work we neglect the role of gas and focus on the other prominent environmental effect in AGNs: the tidal field exerted by the SMBH.
@@
With these motivations, we start in §$\S$2.1 by reviewing the theory of assembling eccentric mergers through binary-single interactions taking place in disc-like environments. We then in §$\S$ 2.2 extend this theory to include the effect from a tidal boundary caused by the presence of the nearby SMBH. After this, we move onto performing PN interactions between a BBH and single BH all orbiting a SMBH (§$\S$ 4.1), for which we present outcome distributions (§$\S$ 4.2 & §$\S$ 4.3) and phase-space diagrams in §$\S$ 4.4. We especially focus on the outcome of probabilities for eccentric mergers as a function of BBH semi-major axis (SMA) relative to the size of the Hill sphere, as well as how the results depend on the inclination angle between the BBH and the incoming single BH (§$\S$ 4.5). Finally, we conclude and highlight future directions in this problem.
@@ 2 Three-body Interactions in a Tidal Field
Figure 1: Illustration of binary-single interaction propagating from left to right resulting in a merger between two of the three objects while the remaining single is still bound (see Sec. 2.1). The system goes through several intermediate states (IMS) before GW capture of two of the three objects. For illustrative purposes, we include horizontal dashed lines to indicate the size of a Hill sphere ($R_{\rm H}$) in case the interaction takes place in the presence of a tidal field. If the interaction extends beyond $R_{H}$, the system becomes unbound and undergoes tidal breakup. The tidal boundary therefore limits the number of interactions $\mathcal{N}$ and it increases the probability of breaking up the system before natural completion. This consequently leads to a reduction in the probability for 3-body and eccentric mergers. The derived correction factor is shown in Eq. 22.
The three-body scattering problem has been studied in great detail with different physical motivations, from the context of classical cluster evolution (Heggie, 1975) to including PN correction terms to follow the formation of GW driven mergers of compact objects (Samsing et al., 2014; Samsing, 2018). While several theoretical studies have been put forward for describing especially the fraction of eccentric merges forming in chaotic 3-body interactions (e.g. Samsing et al., 2014; Rodriguez et al., 2016a; Samsing, 2018), only a few have extended this theory to include how a nearby perturber, such as an SMBH, affects the range, fraction, and nature of outcomes (Leigh et al., 2018; Trani et al., 2019; Ginat & Perets, 2021; Trani et al., 2023; Rom et al., 2023). In S22 it was shown analytically and numerically that co-planar scatterings, such as those that might take place in AGN-disc environments, naturally give rise to an exceptionally high fraction of eccentric GW mergers due to the controlled geometric setup, however neglecting the effect of the SMBH. Below we start by reviewing the relevant 3-body scattering theory, after which we derive how the presence of a nearby object to leading order has an effect on the fraction of eccentric mergers forming during the chaotic state of the 3-body system. This not only adds an important component to this theoretical framework, but also sets the stage and provides motivation for PN scatterings with an SMBH that we present later in the paper.
@@ 2.1 Formation of eccentric mergers
For a 3-body merger that is always highly eccentric at formation, the corresponding GW frequency where most of the power is emitted, also referred as the GW peak frequency (D’Orazio & Samsing, 2018b), is related to the binary pericenter $r_{f}$ distance as,
@@
it consequently shows up eccentric at GW frequency that is higher than $f_{\rm gw}$. It might be that the merger forms with a GW peak frequency that is below the observable band of LVK (Samsing et al., 2014). In this case, one has to evolve the binary until it reaches the band through the emission of GWs; however, at this point the eccentricity might be too low to be measured. Therefore, 3-body mergers are not necessarily representative of the fraction of observable highly eccentric mergers in e.g. LVK as discussed in Samsing (2018). For an IMS merger to clearly show up as an eccentric merger, its GW peak frequency at formation has to be near the observable band.
With the critical distance from Eq. 10, one can easily estimate the probability for the outcome in a similar way to what we did in Eq. 8, which is $\approx\mathcal{N}\times\sqrt{2r_{f}/a_{0}}$. In the more general case for the IMS binary to appear at GW frequency $f_{\rm gw}$ with an eccentricity $e_{f}$, the probability should instead be based on the critical pericenter distance,
@@
as further explained in S22. The question is how these general results, which also have shown excellent agreement with numerical simulations (S22), change when the chaotic three-body interactions take place in a tidal field from a nearby perturber, such as an SMBH. Below we estimate how this addition affects the outcomes.